# cAST: Enhancing Code Retrieval-Augmented Generation with Structural Chunking via Abstract Syntax Tree

### v1

@@ 1 Introduction
Large-scale code generation has emerged as a cornerstone of modern software engineering, powering tasks that range from automated bug fixing (Meng et al., 2024) to full-fledged repository-level completion (Zhang et al., 2023a). Retrieval-augmented generation (RAG) pushes this frontier further by allowing language models to ground their predictions in a rich external corpus of data (Guu et al., 2020), effectively mitigating hallucinations and improving factual correctness (Izacard et al., 2022).
@@
Extensive experiments show that cAST improves performance across a range of code generation tasks. Specifically, it offers three key advantages: (1) Structure-preserving chunks: AST traversal yields more self-contained chunks, improving both retrieval and generation. For instance, StarCoder2-7B sees an average of 5.5 points gain on RepoEval (Zhang et al., 2023b). (2) Cross-language consistency: The language-agnostic nature of cAST enables better generalization across programming languages, achieving up to 4.3 points gain on CrossCodeEval (Ding et al., 2023) (3) Metadata retention: AST-based chunks more faithfully capture metadata at the file, class, and function levels, enhancing context matching in hybrid code+natural language tasks, e.g., up to 2.7 points gain on SWE-bench (Jimenez et al., 2024), which focuses on resolving GitHub issues.
@@
To support syntax-aware chunking, we leverage the Abstract Syntax Tree (AST) representation of code. An AST is a tree-structured abstraction that captures the syntactic structure of source code in a way that is both hierarchical and semantically rich. Rather than treating code as plain text, AST encodes language constructs—like functions, classes, loops, and conditionals—as distinct nodes in a structured parse tree. This enables us to identify meaningful code boundaries with precision, ensuring that chunking respects the underlying syntax. Since ASTs are widely supported across languages, this approach also enhances the language-invariance and portability of our method. Our works uses the tree-sitter library (Tree-sitter, 2025) for the AST tree parsing.
@@ 3 Experiments
We evaluate cAST with various top retrieval and generation models in various code task settings. We present results of selected end-to-end RACG pipelines (retriever + LM) in Section 3.2 and full tables in the Appendix (2, 3, 4, 5).
@@ 3.1 Experiment Settings
Similarly, for generations, we include 2 code-specific LMs: StarCoder2-7B (Lozhkov et al., 2024), CodeLlama-7B-Python (Roziere et al., 2023); and 2 general-purpose ones (claude-3.7-sonnet, gemini-2.5-pro-0325), as both represent the state-of-the-art in coding.
@@
Table 1 shows end-to-end RACG results with selected retrievers (BGE-base, GIST-base, Codesgae-small-v2) on the three datasets. Key observations:
@@
Correlation between retrieval and generation performance
@@
3.3 Related Work
@@
Successful code RAG hinges on pairing high-quality retrievers with generation frameworks that can effectively leverage the fetched context. General-purpose systems—RAG Lewis et al. (2020), FiD Izacard and Grave (2021), and RePlug (Shi et al., 2023)—demonstrate that feeding high-recall evidence to a language model markedly improves factuality. In the software-engineering domain, CodeRAG-Bench (Wang et al., 2024) confirms these gains on repository-level tasks while revealing that lexical-matching retrievers often miss relevant code, motivating code-specific retrieval models. State-of-the-art code retrievers such as CodeBERT Feng et al. (2020), UniXcoder Guo et al. (2022), and CodeRetriever Li et al. (2022) learn joint code–text or code–code embeddings and consistently surpass generic dense models in code search and question answering. Most pipelines still inherit fixed line-based chunking from natural-language RAG. Our work shows that respecting syntactic units with AST-aware chunks further enhances these retrieval-generation loops.
@@ A.6 Ethical Statements
We will open source our code upon acceptance.

### v2

@@ 1 Introduction
Large‑scale code generation has emerged as a cornerstone of modern software engineering, powering tasks that range from automated bug fixing (Meng et al., 2024) to full‑fledged repository-level completion (Zhang et al., 2023a). Retrieval‑augmented generation (RAG) pushes this frontier further by allowing language models to ground their predictions in a rich external corpus of data (Guu et al., 2020), effectively mitigating hallucinations and improving factual correctness (Izacard et al., 2022).
@@
Extensive experiments show that cAST improves performance across a range of code generation tasks. Specifically, it offers three key advantages: (1) Structure-preserving chunks: AST traversal yields more self-contained chunks, improving both retrieval and generation. For instance, StarCoder2-7B sees an average of 5.5 points gain on RepoEval (Zhang et al., 2023b). (2) Cross-language consistency: The language-agnostic nature of cAST enables better generalization across programming languages, achieving up to 4.3 points gain on CrossCodeEval (Ding et al., 2023). (3) Metadata retention: AST-based chunks more faithfully capture metadata at the file, class, and function levels, enhancing context matching in hybrid code+natural language tasks, e.g., up to 2.7 points gain on SWE-bench (Jimenez et al., 2024), which focuses on resolving GitHub issues.
@@
To support syntax-aware chunking, we leverage the Abstract Syntax Tree (AST) representation of code. An AST is a tree-structured abstraction that captures the syntactic structure of source code in a way that is both hierarchical and semantically rich. Rather than treating code as plain text, AST encodes language constructs—like functions, classes, loops, and conditionals—as distinct nodes in a structured parse tree. This enables us to identify meaningful code boundaries with precision, ensuring that chunking respects the underlying syntax. Since ASTs are widely supported across languages, this approach also enhances the language-invariance and portability of our method. Our work uses the tree-sitter library (Tree-sitter, 2025) for the AST tree parsing.
@@ 3 Experiments
We evaluate cAST with various top retrieval and generation models in various code task settings. We present results of selected end-to-end RACG pipelines (retriever + LM) in Section 3.2 and full tables in the Appendix (5, 6, 7, 8).
@@ 3.1 Experiment Settings
Similarly, for generations, we include two code-specific LMs: StarCoder2-7B (Lozhkov et al., 2024), CodeLlama-7B-Python (Roziere et al., 2023); and two general-purpose LMs (claude-3.7-sonnet, gemini-2.5-pro-0325), as both represent the state-of-the-art in coding.
@@
Table 1 presents the end-to-end RACG results with selected retrievers (BGE-base, GIST-base, Codesgae-small-v2) on the three datasets. The results highlight several key observations:
@@
Correlation between retrieval and generation performance.
@@
4 Ablations
Necessity of merging.
The motivation for introducing merging in our algorithm is to maximize the information density of each chunk. Under a split-only approach, small AST nodes, such as import statements and variable assignments, generate an excessive number of chunks, which unnecessarily enlarges the index and degrades retrieval performance. These fine-grained chunks also contain limited context, making them less effective for downstream tasks, as shown in Table 2. Across all retrievers, we find that both retrieval and generation performance decline under the split-only strategy.
| Metric (Model) | Split-then-merge (cAST) | Split-only | |||||
| BGE | GIST | CodeSage | BGE | GIST | CodeSage | ||
| R | nDCG | 71.1 | 75.9 | 85.1 | 53.5 | 59.1 | 66.1 |
| G | Pass@1 (StarCoder2) | 51.7 | 57.9 | 73.2 | 48.3 | 45.0 | 65.4 |
| Pass@1 (CodeLlama) | 49.6 | 56.6 | 72.1 | 47.2 | 48.5 | 58.4 | |
Selection of context length.
In our experiments, we set max_context_length = 4000, which roughly corresponds to the top five chunks. A comparison of different context lengths is shown in Table 3. We observe that doubling the context length does not necessarily improve generation, whereas a modest reduction in context length can lead to performance degradation, likely due to chunk truncation.
| Pipeline (R + G) | Context length (tokens) | |||
| 3500 | 4000 | 8000 | ||
| BGE + StarCoder2 | 46.9 | 51.7 | 51.7 | |
| GIST + StarCoder2 | 57.1 | 57.9 | 58.2 | |
| CodeSage + StarCoder2 | 70.5 | 73.2 | 69.2 | |
Selection of maximum chunk size.
We set max_chunk_size = 2000 in our experiments, as the resulting chunks exhibit similar statistics (e.g., line counts and token counts) to the fixed-size chunking baseline. A sensitivity analysis of max_chunk_size is presented in Table 4. We observe that retrieval and generation performance peak when max_chunk_size is between 2000 and 2500 characters. Additionally, generation performance also depends on max_context_length, as shown in the previous analysis. When context length allows, larger chunks can provide more information, while smaller chunks help mitigate the risk of truncation.
| Metric (Model) | Maximum chunk size | |||||
| 1000 | 1500 | 2000 | 2500 | 3000 | ||
| R | nDCG | 69.0 | 68.4 | 71.1 | 72.3 | 69.4 |
| G | Pass@1 (StarCoder2) | 43.4 | 45.8 | 51.7 | 50.1 | 51.2 |
5 Related Work
@@
Successful code RAG hinges on pairing high-quality retrievers with generation frameworks that can effectively leverage the fetched context. General‐purpose systems—RAG Lewis et al. (2020), FiD Izacard and Grave (2021), and RePlug (Shi et al., 2023)—demonstrate that feeding high-recall evidence to a language model markedly improves factuality. In the software-engineering domain, CodeRAG-Bench (Wang et al., 2024) confirms these gains on repository-level tasks while revealing that lexical-matching retrievers often miss relevant code, motivating code-specific retrieval models. State-of-the-art code retrievers such as CodeBERT Feng et al. (2020), UniXcoder Guo et al. (2022), and CodeRetriever Li et al. (2022) learn joint code–text or code–code embeddings and consistently surpass generic dense models in code search and question answering. Most pipelines still inherit fixed line-based chunking from natural-language RAG. Our work shows that respecting syntactic units with AST-aware chunks further enhances these retrieval-generation loops.
@@ A.6 Ethical Statements
Our code is open source and available at https://github.com/yilinjz/astchunk.